OpenStudy (anonymous):

could someone help to show me the forumula so that i know how much to deposit, bank pays 5%interest compounded semiannually, i need to have 25000 within 6 yrs thanks in advance

OpenStudy (anonymous):

Interest = Principle x Rate x Time

OpenStudy (anonymous):

thx will try that, wish it was for real lol

OpenStudy (anonymous):

homework question

OpenStudy (anonymous):

ok, if i have any more questions then feel free to ask :)

OpenStudy (anonymous):

sure thx

OpenStudy (anonymous):

if calculations right i may have to think about it for real lol

OpenStudy (anonymous):

haha gl getting a bank for that :P

OpenStudy (tkhunny):

This is not simple interest. I = 0.05 -- Annual Interest Rate j = 0.05 / 2 = 0.025 -- Semi Annual Interest Rate \(Deposit * (1 + j)^{6*2} = 25000\) Solve for the Deposit Amount.